OpenStudy (anonymous):

Verify the identity. cos (x - y) - cos (x + y) = 2 sin x sin y

OpenStudy (anonymous):

I got that, what next?

OpenStudy (mathstudent55):

trig identities for sum and difference: cos (a + b) = cos a cos b - sin a sin b cos (a - b) = cos a cos b + sin a sin b cos (x - y) - cos (x + y) = cos x cos y + sin x sin y - (cos x cos y - sin x sin y) = cos x cos y + sin x sin y - cos x cos y + sin x sin y = 2 sin x sin y

OpenStudy (mathstudent55):

That's it. I worked only on the left side, and I turned the left side into the right side, so the identity is proved.

OpenStudy (anonymous):

Oh!

OpenStudy (anonymous):

Thanks!

OpenStudy (mathstudent55):

wlcm